# Noemi Walczak Université Paris 1 Panthéon-Sorbonne, Paris, France Kenan Huremović Armando Rungi

### v3

@@ 4. QUANTIFYING THE EFFECT OF THE CBAM
We use exact hat-algebra as developed by [13] to characterize a counterfactual equilibrium in terms of proportional changes relative to the steady state. Specifically, the equilibrium in relative changes is computed by solving the following system of non-linear equations that jointly determine changes in wages, intermediate input prices, carbon prices and emissions, cost and consumption shares, expenditures, and accounts for the endogenous nature of the CBAM.
@@
    $\begin{cases}\hat{t}_{i}=\frac{1}{t_{i}\overline{E_{i}}}\sum_{j\in\mathcal{J}}\rho_{i}^{j}p_{i}^{j^{\prime}}q_{i}^{j^{\prime}}\quad\forall i\in\mathcal{N}_{c}\\ \hat{E}_{i}=\frac{1}{\overline{t_{i}}E_{i}}\sum_{j\in\mathcal{J}}\rho_{i}^{j}p_{i}^{j^{\prime}}q_{i}^{j^{\prime}}\quad\forall i\in\mathcal{N}_{nc}\end{cases}$ (16)
The endogenous nature of CBAM is visible from equation (10), where the part of the total tariffs which is due to CBAM ($\rho_{n}^{k}\frac{t^{\prime}_{i}}{t^{\prime}_{n}}$) depends on the counterfactual values of prices. In the case of exogenous CBAM, this term changes to ($\rho_{n}^{k}\frac{t_{i}}{t_{n}}$), where $t_{i}$ and $t_{n}$ are the price levels in the equilibrium prior to the introduction of CBAM.
We solve the system using a fixed point iteration algorithm, similar to the one used in [6, 24], following the steps below:
@@
    Guess the initial vectors of changes for $\mathbf{w}$ and $\mathbf{t}$;
@@
Solving the model in relative changes has the advantage of reducing the data needed for the calibration. We rely on a minimal yet sufficient set of macroeconomic and environmental data, which enhances the transparency, tractability, and ease of replication of our exercise. Specifically, we calibrate the model on data for the last available year 2018888See A.1 for a detailed exposition of the mapping of the model to the data. Following [3], we construct the baseline year (2024) by incorporating environmental policies implemented between 2018 and 2024. The effects of the CBAM are then assessed relative to this baseline. Further details are provided later in this section.. In particular, we work on bilateral trade flows in intermediate inputs and final goods, gross output and value added sourced from the OECD Inter-Country Input-Output tables [10], which collects data from 44 sectors across 32 countries, complemented by a residual Rest of the World. With these data, we can retrieve the empirical counterparts for the steady state values of the cost shares $\iota_{ni}^{kj}$, the labor input shares $\beta_{i}^{j}$, the consumption shares $\chi_{i}^{j}$ and the exogenous deficits $\bar{D_{i}}$. Then, we source tariffs from the WTO Integrated Database and Consolidated Tariff Schedule [^9] We follow [2], and calibrate the model using observed trade value data that we adjust for existing tariffs. This practically means that in our quantitative exercise, we set $\kappa_{ni}^{kj}=0$ at the baseline., while we retrieve Scope 1 emissions from production from the OECD Environmental Statistics database. In order to build the emissions elasticity $\rho_{i}^{j}$, emissions quantities are multiplied by the corresponding country’s Effective Carbon Rate [26], converted into USD using the OECD exchange rate. For the European Union, Iceland and Norway, we employ a different approach. We collect emissions data from the European Union Transaction Log (EUTL) database. We aggregate these data at the selected sectoral level and adjust the reported emissions by deducting those covered by free allowances granted to the installations operating in Energy Intensive Trade Exposed (EITE) industries. Note that this choice implies that we will evaluate lower bounds for both the level and the variation of emissions in these countries, as the data source excludes emissions from entities not covered by the ETS. For this reason, we refer to EU emissions and ETS emissions interchangeably, unless the distinction is necessary to avoid ambiguity.
@@
In countries with an operational national carbon pricing scheme by 2024, we set the total supply of emissions to match the aggregate quantities we retrieve from the respective databases, described above. For countries without a pricing scheme in place by 2024101010We collect details about direct carbon pricing initiatives around the world on the World Bank Carbon Pricing Dashboard (https://carbonpricingdashboard.worldbank.org/). The full list includes: Bangladesh, Brazil, Côte d’Ivoire, Costa Rica, Egypt, India, Israel, Morocco, Malaysia, Nigeria, Peru, Philippines, Russia, Turkey, and the United States. Moreover, we include China in this list, since we believe its current environmental regulation is not stringent enough to incentivize decarbonisation efforts., we impose an exogenous carbon price equal to the country-specific Effective Carbon Rate, while the supply of emissions is determined endogenously. Finally, consistent with [7], we assume an elasticity of substitution between intermediate goods and final consumption goods equal to 4, constant across countries and sectors.
@@
Before introducing the CBAM, we run a simulation in which, following [3], selected countries implement their emissions reduction pledges under the Paris Agreement. Then we use the resulting equilibrium as our baseline. Rather than comparing the scenario where the CBAM is introduced with the original 2018 steady state, we adopt this policy-adjusted baseline for two main reasons. First, the implementation of the Paris Agreement commitments is binding for many countries, making it more representative of the environment where the CBAM operates. Second, it allows us to isolate the specific contribution of the CBAM, net of the ongoing climate policy changes.
We build the policy-adjusted baseline in the following way. We select the countries with a carbon pricing scheme and consider the unconditional targets declared in their 2020 Nationally Determined Contributions (NDCs) under the Paris Agreement [^11] We consider the most recent unconditional pledges submitted to the NDCs Registry as of 2024. Conditional targets are excluded from this exercise.. We convert all the considered commitments into a fixed target level for the year 2030 and derive a yearly percentage reduction rate from the declared base year, consistent with each country’s target. We then apply an exponential decay from 2018 to 2024, in order to model emission trajectories. Two notable exceptions apply. First, as for ETS countries, we also consider a 40% reduction in the freely allocated allowances relative to 2018 levels, as supported by empirical data (see Figure A1). Second, as for China, we follow [3] in assuming the absence of a fully operational carbon market. We believe that its current carbon price is too low to reach the targets declared in the country’s NDC [^12] According to the [26], in the last year recorded, coinciding with 2023, the Effective Carbon Rate was EUR 7.27 per tonne of $\text{CO}_{2}\text{e}$. and, accordingly, we treat China’s carbon price as exogenous.
We impute the computed change in emissions between 2018 and 2024 and simulate the counterfactual equilibrium. The resulting cost-shares and consumption shares, together with the new pollution intensity parameters, are then used to recalibrate the model and build our new baseline.
@@
  • Current CBAM: under the Regulation, the border adjustment initially applies only to specific goods – iron and steel, cement, aluminum, organic basic chemicals, hydrogen, fertilizers, and electricity. In our exercise, they correspond to six broader sectors: Mining and Quarrying, Chemicals and Chemical Products, Non-Metallic Mineral Products, Basic Metals, Fabricated Metal Products, and Electricity, Gas, Steam, and Air Conditioning Supply;
@@ 5. RESULTS
In this section, we present our main findings. We begin by analyzing how the CBAM affects international trade patterns, focusing on the following three outcomes: the average share of intermediate inputs the EU imports from abroad, the share of intermediate inputs EU sources domestically, and the size of EU sectors as measured by their average Domar weight in the world economy. We report these results in Table 1. Next, we turn to our core findings on carbon leakage and emissions embodied in EU imports, summarized in Table 2. We conclude by evaluating the impact of the CBAM on gross national expenditure in both the EU and non-EU countries, and report the results in Table 3. For each outcome, we compare the implications of the current versus full CBAM and distinguish between scenarios in which the value of the CBAM is set endogenously and exogenously, as discussed in Section 3. Our preferred specification is the one with current and endogenous CBAM, reported in the first column of Tables 13.
@@
When discussing our results in the following paragraphs, we will use the terms dirty/carbon-intensive goods to refer to the sectoral goods falling under the ETS scheme [^13] See Table A1 for the classification of sectors falling under the ETS., while classifying the residual sectoral goods as clean/non-carbon-intensive.
@@
We start by examining how the patterns of international trade are affected by the CBAM. By design, the CBAM increases the price of imported carbon-intensive goods at the EU border, creating a price wedge that alters the relative competitiveness of domestic and foreign producers. Therefore, we first ask a question about whether and to what extent the introduction of the CBAM causes the European firms to substitute away from dirty imported inputs.
As reported in Panel (a) in Table 1, this mechanism leads, under the current CBAM, to a decline in the average share of foreign purchases – corresponding to the average share of European imports from all other countries in total intermediate input purchases – of 0.60%. This number more than doubles in the full CBAM scenario. The reported effects are similar across endogenous and exogenous CBAM scenarios. The additional adjustment of the value of CBAM in the endogenous scenario leads to an increase of the effect by 5% (from 0.57% to 0.60%) under the current CBAM, and 3% under full CBAM. At a more granular level, while the CBAM leads to a contraction in the shares of foreign purchases of carbon-intensive sectoral goods directly targeted by the policy (with a 1.06% decrease under the current CBAM, reaching 2.14% under the full CBAM), there is a reallocation of the input demand towards cleaner alternatives driven by substitution effects, with an increase in the average share of foreign purchases of non-carbon-intensive goods of 0.57% under the current CBAM scenario, expanding to 1.91% under the full CBAM. [^14] The magnitude of these substitution effects clearly depends on the elasticity of substitution across intermediate inputs, captured with $\theta$. As shown in Table A3, a higher elasticity amplifies the reallocation from dirty to cleaner inputs, magnifying the effects of the policy.
Panel (b) offers a complementary perspective to Panel (a) and presents the effect on the share of intermediate inputs sourced within the EU. The additional cost introduced by the CBAM leads to a partial substitution away from more expensive foreign intermediate inputs in favour of domestically produced goods, inducing a reallocation of purchases of intermediate goods towards the EU, more so for clean than for dirty inputs. Under the current CBAM, domestic purchases of clean goods rise by 0.08%, while those of polluting goods increase by 0.03%. These effects are widened under the full CBAM, with domestic purchases of clean goods rising to 0.10% and those of polluting goods to 0.07%.
Finally, Panel (c) in Table 1 shows the average change in Domar weights (defined in Section 3). Despite the small albeit positive overall change, disaggregated results reveal a difference between clean and dirty goods. The clean sectors become bigger, since other sectors substitute dirty foreign inputs with domestic clean inputs. The average Domar weight of clean sectors exhibits an increase of 0.11% under the current CBAM (0.29% under the full CBAM), whereas the corresponding share of dirty goods declines marginally in both scenarios. This asymmetric effect stems from the different reliance of clean and dirty producers on carbon-intensive inputs: while the CBAM raises unit production costs for all producers by increasing the price index, clean producers are less exposed to this shock due to their lower input carbon intensity. As a result, their relative competitiveness increases, while European carbon-intensive sectors suffer a loss of sales shares due to a shift in the geographical sourcing of such goods. In fact, importers in other countries are more likely to reallocate their demand away from the European Union and towards third-country producers.
The comparison of the cases with endogenous vs. exogenous CBAM shows that the substitution effects in intermediate inputs are magnified by the endogenous adjustment, as a result of the feedback loop between the CBAM and the equilibrium carbon prices. As the CBAM increases the relative price of foreign carbon-intensive goods, demand shifts towards cleaner and domestic substitutes. This reallocation reduces production abroad and raises it at home, leading to a decrease in foreign carbon prices – due to lower carbon input demand – and an increase in the domestic carbon prices. Since the CBAM responds negatively to foreign carbon prices and positively to domestic ones, the differential widens, increasing the value of the CBAM. These general equilibrium feedbacks magnify the substitution away from foreign inputs. Hence, the magnification in Panel (a) when moving from exogenous towards endogenous CBAM.
Variable Current CBAM Full CBAM
@@
Panel (a) Average $\Delta\%$ in EU Share of Foreign Purchases
Total -0.60 (1.30)
-0.57
-1.27 (2.80)
-1.23
   …of clean intermediate goods 0.57 (0.34)
0.55
1.91 (1.70)
1.86
   …of dirty intermediate goods -1.06 (2.37)
-1.02
-2.14 (3.59)
-2.07
Panel (b) Average $\Delta\%$ in EU Share of Domestic Purchases
@@
Panel (c) Average $\Delta\%$ in EU Domar Weights
Total 0.03 (0.19)
0.03
0.14 (0.40)
0.14
   …of clean intermediate goods 0.11 (0.09)
0.10
0.29 (0.29)
0.28
   …of dirty intermediate goods -0.03 (0.22)
-0.02
0 (0.44)
0
@@
Table 1: Policy-induced changes in EU-ETS shares (% of baseline). Panel (a) presents the mean percentage change across countries of the share EU purchases of intermediate inputs sourced from non-EU countries in EU total intermediate input purchases, with standard deviations (in parentheses) capturing the dispersion of country-specific changes around the mean. Panel (b) shows the mean percentage change of the share of intermediate inputs sourced domestically. Finally, Panel (c) reports the mean percentage change in EU sectoral sales, with the standard deviations (in parentheses) indicating the dispersion of sectoral changes around this average. Each value is considered under two scenarios: one in which the CBAM is determined endogenously (columns labeled Endog.) and one in which it is exogenous (columns labeled Exog.). Our main analysis focuses on the endogenous scenario (columns Endog.).
The effect of CBAM is heterogeneous across trade partners. The magnitude of the border adjustment varies across countries and sectors due to three key factors: (i) the pollution intensity of the production process; (ii) the relative stringency of carbon pricing; (iii) the sectoral composition of exports to the EU. These heterogeneous effects are illustrated by Figure 2, which depicts the change in the share of carbon-intensive foreign purchases for the ten largest exporting countries of polluting inputs to the EU (Panel 2(a)). Interestingly, the introduction of the CBAM leads to an increase in shares of purchases from some of these countries. The increase in the share from Switzerland and Norway is straightforward to interpret: since the CBAM does not apply to imports from countries that are either part of the EU-ETS or have a domestic ETS fully linked with it (as in the case of Switzerland), their exports are not subject to the additional carbon price wedge. As a result, these countries gain a relative cost advantage compared to exporters from countries whose goods are targeted by the CBAM, which translates into increased export shares to the EU. The cases of Japan and South Korea are different. Both countries are major exporters of carbon-intensive goods to the EU, yet they display distinctive patterns in response to the CBAM with respect to the other trading partners. This is because Japan has relatively low carbon intensity in the production of carbon-intensive goods (see Table A2), while South Korea has a relatively high carbon price (see Figure A4). Both factors lead to low CBAM relative to these countries. When we consider the reduction in the shares of foreign purchases from the top 10 most affected countries (Panel 2(b)) our results are aligned with the Relative CBAM Exposure Index [^15] https://www.worldbank.org/en/data/interactive/2023/06/15/relative-cbam-exposure-index, with Ukraine being the most affected country, followed by South Africa, India, Russia, and China. All countries that are identified as highly exposed to the CBAM in the index experience a decline in their export share towards the European Union.
Figure 3 illustrates the heterogeneous impact of the CBAM on the Domar weights of European sectors, highlighting the 15 most affected industries across both carbon-intensive and non-carbon-intensive categories. The most substantial declines occur in carbon-intensive sectors — such as basic metals, chemical products, and fabricated metal products. In contrast, non-carbon-intensive sectors, which are not directly targeted by the current CBAM, experience marginal or modest gains. These gains arise from the complex substitution patterns along the supply chain and are captured by our model. As an illustration, consider Coke and refined petroleum products sector. When only a subset of sectors is targeted under the current CBAM, Coke and refined petroleum products, that is not subject to the wedge, act as a substitute for regulated inputs, experiencing an increase in its size. However, when the CBAM is extended to cover all sectors under the EU-ETS, including Coke and refined petroleum products, the effect is reversed and the competitive advantage of the sector vanishes. These indirect effects become increasingly significant as the elasticity of substitution increases, as shown in sensitivity checks in Appendix Figure A6.
@@
Figure 2: Policy-induced changes in European shares of foreign purchases of dirty goods for selected countries (% of baseline level).
Figure 3: Policy-induced changes in European Domar weights of selected sectors (% of baseline level).
We now turn to the impact of CBAM on emissions, beginning with the emissions embodied in imports.
Emissions embodied in trade are calculated using the vector of production-based emissions and input-output multipliers. We borrow the definition from the OECD [30] and we adapt it to align with the structure of our model. The emissions embodied in European imports are given by the following vector $\mathbf{EEI}(EU)\in\mathbb{R}^{NJ}$ with entries corresponding to the emissions generated from the production of each good $j\in\mathcal{J}$ in every country $i\in\mathcal{N}$ embodied in European imports. In particular, the first equivalence in equation (17) corresponds to the OECD definition, while the subsequent equality restates it in terms of our notation and modeling framework [^16] A more detailed derivation is left to the corresponding section in the Appendix A.2.:

### v4

@@ 5. Quantifying the effect of CBAM
We use the exact hat-algebra following [dekle2008global] to characterize a counterfactual equilibrium in terms of proportional changes relative to the steady state. Specifically, the equilibrium in relative changes is computed by solving the following system of nonlinear equations that jointly determine changes in wages, intermediate input prices, carbon prices, and emissions, as well as cost and consumption shares, and expenditures.
@@
    $\begin{cases}\hat{t}_{i}=\frac{1}{t_{i}\overline{E_{i}}}\sum_{j\in\mathcal{J}}\rho_{i}^{j}p_{i}^{j^{\prime}}q_{i}^{j^{\prime}}\qquad\forall i\in\mathcal{N}_{c}\\ \hat{E}_{i}=\frac{1}{\overline{t_{i}}E_{i}}\sum_{j\in\mathcal{J}}\rho_{i}^{j}p_{i}^{j^{\prime}}q_{i}^{j^{\prime}}\qquad\forall i\in\mathcal{N}_{nc}\end{cases}$ (16)
Please note that the endogenous nature of CBAM is visible from equation (10), where the part of the total tariffs which is due to CBAM ($\rho_{n}^{k}\frac{t^{\prime}_{i}}{t^{\prime}_{n}}$) depends on the counterfactual values of prices. In the case of exogenous CBAM, this term changes to ($\rho_{n}^{k}\frac{t_{i}}{t_{n}}$), where $t_{i}$ and $t_{n}$ are the price levels in the equilibrium prior to the introduction of CBAM.
We solve the system using a fixed-point iteration algorithm, similar to the one used in [caliendo2015estimates, mayer2019cost], following the steps below:
@@
    Guess the initial vectors of changes for $\mathbf{w}$, $\mathbf{t}\in\mathbb{R}^{N}$;
@@
Solving the model in relative changes has the advantage of reducing the data needed for the calibration. We rely on a minimal yet sufficient set of macroeconomic and environmental data, which enhances the transparency, tractability, and ease of replication of our exercise. Specifically, we calibrate the model using data from the last available year, 2018 (see A.1 for a detailed exposition of the model’s mapping to the data). Following [bellora2023eu], we construct the baseline year (2024) by incorporating environmental policies implemented between 2018 and 2024. This is important to account for the reduction in energy-related $\text{CO}_{2}$ emissions observed in advanced economies during this period. The effects of the CBAM are then assessed relative to the baseline year 2024. Further details are provided later in this section.
@@
In particular, we work on bilateral trade flows in intermediate inputs and final goods, gross output and value added sourced from the OECD Inter-Country Input-Output tables [oecd2021], which collect data from 44 sectors across 32 countries, complemented by a residual Rest of the World to close the global matrix. With these data, we can retrieve the empirical counterparts for the steady state values of the cost shares $\iota_{ni}^{kj}$, the labor input shares $\beta_{i}^{j}$, the consumption shares $\chi_{i}^{j}$, and the exogenous deficits $\bar{D_{i}}$. Then, we source tariffs from the WTO Integrated Database and Consolidated Tariff Schedule [^9] We follow [baqaee2024networks] and calibrate the model using observed trade value data and emissions, adjusting for existing tariffs. This practically means that we set $\kappa_{ni}^{kj}=0$ at the baseline in our quantitative exercise., while we retrieve Scope 1 emissions from production from the OECD Environmental Statistics database. In order to build the emissions’ elasticity $\rho_{i}^{j}$, emissions quantities are multiplied by the corresponding country’s Effective Carbon Rate [ECR2023], converted into USD using the OECD exchange rates. For the European Union, Iceland, and Norway, we employ a different approach. We collect emissions data from the European Union Transaction Log (EUTL) database. We aggregate them at the selected sectoral level and adjust the reported emissions by deducting those covered by free allowances granted to the installations operating in Energy Intensive Trade Exposed (EITE) industries. Note that this choice implies that we will evaluate lower bounds for both the level and the variation of emissions in these countries, as the data source excludes emissions from entities not covered by the ETS. For this reason, we refer to EU emissions and ETS emissions interchangeably, unless the distinction is necessary to avoid ambiguity.
@@
In countries with an operational national carbon pricing scheme by 2024, we set the total supply of emissions to match the aggregate quantities we retrieve from the databases described above. For countries without a pricing scheme in place by 2024101010We collect details about direct carbon pricing initiatives around the world on the World Bank Carbon Pricing Dashboard (https://carbonpricingdashboard.worldbank.org/). The full list includes: Bangladesh, Brazil, Côte d’Ivoire, Costa Rica, Egypt, India, Israel, Morocco, Malaysia, Nigeria, Peru, Philippines, Russia, Turkey, and the United States. Moreover, we include China in this list, since we believe its current environmental regulations are not stringent enough to incentivize decarbonisation efforts., we impose an exogenous carbon price equal to the country-specific Effective Carbon Rate, while the supply of emissions is determined endogenously. Finally, consistent with [caliendo2022distortions], we assume an elasticity of substitution between intermediate goods and final consumption goods equal to 4, constant across countries and sectors.
@@
Before introducing the CBAM, we run a simulation in which, following [bellora2023eu], selected countries implement their emissions reduction pledges under the Paris Agreement. Then, we use the resulting equilibrium as our baseline. Rather than comparing the scenario where the CBAM is introduced with the original 2018 steady state, we adopt this policy-adjusted baseline for two main reasons. First, the implementation of the Paris Agreement commitments is binding for many countries, making it more representative of the environment where the CBAM operates. Second, it enables us to isolate the specific contribution of the CBAM, net of ongoing climate policy changes and the related adjustments in emissions observed in recent years.
For the baseline, we select the countries with a carbon pricing scheme and consider the unconditional targets declared in their 2020 Nationally Determined Contributions (NDCs) under the Paris Agreement [^11] We consider the most recent unconditional pledges submitted to the NDCs Registry as of 2024. Conditional targets are excluded from this exercise.. We convert all the considered commitments into a fixed target level for the year 2030 and derive a yearly percentage reduction rate from the declared base year, consistent with each country’s target. We then apply an exponential decay from 2018 to 2024 to model emission trajectories. Two notable exceptions apply. First, regarding ETS countries, we also consider a 40% reduction in the freely allocated allowances relative to 2018 levels, as supported by empirical data (see Figure A1). Second, as for China, we follow [bellora2023eu] in assuming the absence of a fully operational carbon market. We believe that its current carbon price is too low to reach the targets declared in the country’s NDC [^12] According to the [ECR2023], in the last year recorded, coinciding with 2023, the Effective Carbon Rate was EUR 7.27 per tonne of $\text{CO}_{2}\text{e}$. and, accordingly, we treat China’s carbon price as exogenous.
We impute the computed change in emissions between 2018 and 2024 and simulate the counterfactual equilibrium. The resulting cost shares and consumption shares, together with the new pollution intensity parameters, are then used to recalibrate the model and build our new baseline.
@@
  • Reduced CBAM: under the Regulation, the border adjustment initially applies only to specific goods – iron and steel, cement, aluminum, organic basic chemicals, hydrogen, fertilizers, and electricity. In our exercise, they correspond to six broader sectors: Mining and Quarrying, Chemicals and Chemical Products, Non-Metallic Mineral Products, Basic Metals, Fabricated Metal Products, and Electricity, Gas, Steam, and Air Conditioning Supply;
@@ 6. Results
In the following paragraphs, we present our main findings. We begin by analyzing how the CBAM affects international trade patterns, focusing on the following three outcomes: the average share of intermediate inputs the EU imports from abroad, the share of intermediate inputs the EU sources domestically, and the size of EU sectors as measured by their average Domar weights in the world economy. We report these results in Table 1. Next, we turn to our core findings on carbon leakage and emissions embodied in EU imports, summarized in Table 2. We continue by evaluating the impact of the CBAM on welfare gains, as proxied by gross national expenditure (GNE) in both the EU and extra-EU countries in Table 3. For each outcome, we compare the implications in our $2\times 2$ scenarios. On the one hand, we present the reduced versus full CBAM and, on the other hand, we distinguish between scenarios in which the value of the CBAM is set either endogenously or exogenously, as discussed in Section 4.
@@
Our preferred specification is the one with full and endogenous CBAM, which corresponds to the scenario in which the original policy design is fully in effect. When discussing our results in the following paragraphs, we will use the terms dirty/carbon-intensive goods to refer to the sectoral goods falling under the ETS scheme [^13] See Appendix Table A1 for the classification of sectors falling under the ETS and/or the CBAM., while classifying the residual sectoral goods as clean/non-carbon-intensive.
Finally, we conclude our contribution with two counterfactual exercises to separate the roles of international trade and technology in Figure 5. Ceteris paribus, we first vary the output elasticity of emissions, then the share of imported inputs.
6.1 Trade patterns.
@@
Let us start with trade patterns. By design, the CBAM immediately increases the relative price of imported carbon-intensive goods at the EU border, by creating a price wedge that worsens the relative competitiveness of foreign producers. Therefore, we first check to what extent CBAM causes substitution away from dirty imported inputs. In panel (a) of Table 1, this mechanism leads to a decline in the average share of foreign purchases, which is the average share of European imports from all other countries in total intermediate input purchases. In case of a reduced CBAM, the change in the foreign purchases share is $0.6\%$, while in the case of the full CBAM it more than doubles to $1.27\%$.
Variable Reduced CBAM Full CBAM
@@
Mean St.dev. Mean St.dev. Mean St.dev. Mean St.dev.
Panel (a) $\Delta\%$ in EU Share of Foreign Purchases
Total -0.60 1.30 -0.57 1.25 -1.27 2.80 -1.23 2.73
   …of clean intermediate goods 0.57 0.34 0.55 0.33 1.91 1.70 1.86 1.66
   …of dirty intermediate goods -1.06 2.37 -1.02 2.28 -2.14 3.59 -2.07 3.49
Panel (b) $\Delta\%$ in EU Share of Domestic Purchases
@@
Panel (c) $\Delta\%$ in EU Domar Weights
Total 0.03 0.19 0.03 0.19 0.14 0.40 0.14 0.39
   …of clean intermediate goods 0.11 0.09 0.10 0.09 0.29 0.29 0.28 0.28
   …of dirty intermediate goods -0.03 0.22 -0.02 0.22 0 0.44 0 0.43
@@
Table 1: Panel (a) presents the average percentage change across countries of the share that the EU purchases of intermediate inputs sourced from non-EU countries in EU total intermediate input purchases (% of baseline year 2024). The standard deviations indicate heterogeneity of country-industry observations. Panel (b) shows the average percentage change of the share of intermediate inputs sourced domestically. Finally, Panel (c) reports the average percentage change in EU sectoral sales, with the corresponding standard deviations. Each value is considered under different scenarios: a full and a reduced version of CBAM can be combined with endogenous (Endog.) or exogenous (Exog.) prices. The original design of CBAM is the full version with endogenous prices.
Please note how reported effects are not that different across the endogenous and exogenous CBAM scenarios; that is, the additional adjustment in the endogenous scenario leads to a mere 5% increase in the effect (from $0.57\%$ to $0.60\%$) under reduced CBAM, and a 3% increase (from $1.23\%$ to $1.27\%$) under full CBAM. Based on this evidence, one might think that the effort to design a policy that is endogenous to carbon pricing is futile. Effects will be similar across all dimensions in the following paragraphs, when we show results on emissions and gains from trade. Yet, a caveat is needed. Our simulations are based on what happens in the current carbon market, in the absence of that strategic element that is crucial in the CBAM policy design. In fact, we have not yet observed any changes in technology by foreign producers in response to the actual CBAM, and there has been no establishment of additional national carbon markets by foreign partners. Both responses are foreseen in the original policy design. In this case, we believe our results for endogenous scenarios remain valid, and they can be considered the lower bound of what would happen in the absence of strategic responses.
Let us describe how the feedback loop between the CBAM and equilibrium carbon prices operates in the current institutional environment. First, CBAM increases the relative price of foreign carbon-intensive goods; thus, demand shifts towards cleaner and domestic substitutes. This reallocation reduces production abroad and increases it at home, resulting in a decrease in foreign carbon prices due to lower demand for carbon-intensive inputs. Since the CBAM responds negatively to foreign carbon prices and positively to domestic ones, the differential widens, increasing the value of the CBAM. This general-equilibrium feedback magnifies the substitution away from foreign inputs. Hence, the albeit small magnification in panel (a), when moving from exogenous toward endogenous CBAM.
At a more granular level, while the CBAM leads to a contraction in the shares of foreign purchases of carbon-intensive sectoral goods directly targeted by the policy (with a 1.06% decrease under the reduced CBAM, reaching 2.14% under the full CBAM), there is a reallocation of the input demand towards cleaner alternatives driven by substitution effects, with an increase in the average share of foreign purchases of cleaner goods by 0.57% under the reduced CBAM scenario, expanding to 1.91% under the full CBAM. Please note how the magnitude of these substitution effects clearly depends on the elasticity of substitution across intermediate inputs, captured with $\theta$. As shown in Appendix Table A3, a higher elasticity amplifies the reallocation from dirty to cleaner inputs.
Panel (b) offers a complementary perspective to Panel (a). It presents the effect on the share of intermediate inputs sourced within the EU. The additional cost introduced by the CBAM leads to a partial substitution away from more expensive foreign intermediate inputs in favour of domestically produced goods, more so for clean than for dirty inputs. Under the reduced CBAM, domestic purchases of clean goods rise by 0.08%, while those of polluting goods increase by 0.03%. These effects widen under the full CBAM, with domestic purchases of clean goods rising to 0.18% and those of polluting goods to 0.07%.
Finally, panel (c) in Table 1 shows the average change in EU Domar weights (defined in Section 4). This is necessary to understand how the size of the EU economy changes. Despite the small albeit positive overall change, disaggregated results reveal a difference between clean and dirty goods. The clean sectors slightly expand, and the dirty sectors slightly shrink. The average Domar weight of clean sectors exhibits an increase of 0.11% under the reduced CBAM (0.29% under the full CBAM), whereas the corresponding share of dirty goods declines marginally under the reduced CBAM and remains unchanged under the full CBAM. This asymmetric effect stems from the different reliance of clean and dirty producers on carbon-intensive inputs: while the CBAM raises unit production costs for all producers, clean producers are less exposed to this shock due to their lower input carbon intensity. As a result, their relative competitiveness increases, while European carbon-intensive sectors suffer a loss of sales. In fact, importers in other countries are more likely to reallocate their demand away from the European Union and towards third-country producers. Overall, the total EU Domar weights slightly increase, thus indicating a modest growth of the EU contribution to global production networks.
Notably, the effect of CBAM is heterogeneous across trade partners. This is shown by the standard deviations of Table 1. The magnitude of the border adjustment varies across countries and sectors due to three key factors: (i) the carbon intensity of the production process; (ii) the relative stringency of carbon pricing; (iii) the sectoral composition of exports to the EU. These heterogeneous effects are illustrated in Figure 2, which shows the change in the share of carbon-intensive foreign purchases for the ten largest exporting countries of polluting inputs to the EU (panel 2(a))  [^14] See Figure LABEL:share_EU_imports_appendix for the baseline shares of EU Carbon Intensive Imports by country.. Interestingly, the introduction of the CBAM leads to an increase in shares of purchases from some of these countries.
@@
Figure 2: Changes in European shares of foreign purchases of dirty goods for selected countries (% of baseline year 2024).
The increase in the share from Switzerland and Norway is straightforward to interpret: since the CBAM does not apply to imports from countries that are part of the EU-ETS, although they are not members of the EU (as in the case of Switzerland), their exports are not subject to the additional carbon price wedge. Briefly, producers located in non-EU countries within the European Economic Area benefit from a relative cost advantage compared to exporters from countries targeted by the CBAM, which translates into increased export shares to the EU. The latter is a significant side effect of CBAM that has been largely overlooked so far.
The cases of Japan and South Korea are different. Both countries are major exporters of carbon-intensive goods to the EU, yet they exhibit distinct patterns in their trade relations with other partners. This is because Japan has a relatively low carbon intensity in the production of carbon-intensive goods (see Appendix Table A2), while South Korea has a relatively high carbon price (see Appendix Figure A4). Both factors lead to low CBAM relative to these countries. On the other hand, when we consider changes in the shares of foreign purchases, Ukraine is the most affected country, followed by South Africa, India, Russia, and China [^15] Please note how our results align with the Relative CBAM Exposure Index reported by the World Bank. All countries identified as highly exposed to the CBAM in the index experience a decline in their export share to the European Union..
Figure 3: Changes in European Domar weights of selected sectors (% of baseline year 2024).
Finally, Figure 3 illustrates the heterogeneous impact of CBAM on the EU Domar weights, highlighting the 15 most affected industries across both carbon-intensive and non-carbon-intensive categories. The most substantial declines occur in carbon-intensive sectors — such as basic metals, chemical products, and fabricated metal products. In contrast, non-carbon-intensive sectors, which are not directly targeted by the reduced CBAM, experience marginal or modest gains. Such gains arise from substitution patterns along the supply chain, as captured by our model. As an illustration, consider Coke and refined petroleum products. When only a subset of sectors is targeted under the reduced CBAM, Coke and refined petroleum products, that is not subject to the wedge, act as a substitute for regulated inputs, experiencing an increase in its size. However, when the CBAM is extended to cover all sectors under the EU-ETS, including Coke and refined petroleum products, the effect is reversed and the competitive advantage of the latter sector vanishes. These indirect effects become increasingly significant as the elasticity of substitution increases, as shown in sensitivity checks in Appendix Figure A6.
6.2 Emissions.
We start by adapting the definition of emissions from the OECD [yamano2020co2] to our theoretical framework. Let us consider the following equation, which indicates the emission embodied in European imports: